I need to write a function given these four zeros
-1
2
3+i
3-1

Answer by stanbon(75887)

I need to write a function given these four zeros
-1
2
3+i
3-1
----------------
f(x) = (x+1)(x-2)((x-3)-i)((x-3)+i)
---
f(x) = (x+1)(x-2)((x-3)^2+1)
---
f(x) = (x+1)(x-2)(x^2-6x+10)
==================================
Cheers,
Stan H.